OpenStudy (anonymous):

The first five terms of a sequence are shown below. 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? A. f(n) = 5n - 1; n ≥ 1 B. f(n) = 5n + 1; n ≥ 1 C. f(n) = 4n - 1; n ≥ 1 D. f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

@Ethanxiaorox

OpenStudy (solomonzelman):

Do you know the difference?

OpenStudy (anonymous):

@roz_mar23

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Y cant u just use +4....... :P

OpenStudy (anonymous):

Solomon i think he is just looking for the answer

OpenStudy (anonymous):

but we should help him

OpenStudy (solomonzelman):

find the difference, that will be the "slope" of it.

OpenStudy (anonymous):

5+(n-1)4=4n+1

OpenStudy (anonymous):

n>=1

OpenStudy (anonymous):

No I do not, I got c earlier

OpenStudy (solomonzelman):

\(\Large \color{black}{ d=a_{n}~-~a_{(n-1)} }\)

OpenStudy (anonymous):

first term=a=5 common difference=d=4 formula of nth term=a+(n-1)d=5+(n-1)4=4n+1:n >=1

OpenStudy (anonymous):

@IndianBoyNiko

OpenStudy (anonymous):

answer ?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

If f(x) = 7x - 1, what does f(12) represent? A. The value if x when (7x - 1) = 12 B. The value of (7x - 1) when x = 12 C. The value of 12 f(x) when x = 12 D. The value of f(x) when (7x - 1) = 12

OpenStudy (anonymous):

why do you ask this question??? The answer is B

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

The value of (7x - 1) when x = 12 option B

OpenStudy (anonymous):

The graphs of functions f(x) and g(x) = f(x) + k are shown below:

OpenStudy (anonymous):

The value of k is ___.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

i'm bad at graphs because i'm stupid